OpenStudy (anonymous):

completing a square: 3a^2+18a+76=8

zepdrix (zepdrix):

Hey tiff, \(\Large\bf \color{#CC0033}{\text{Welcome to OpenStudy! :)}}\)

zepdrix (zepdrix):

Completing the Square huh? Mmmm let's see..

zepdrix (zepdrix):

We would like our equation to be in this form, \[\Large\rm \color{orangered}{x^2+bx}=stuff\]So we can complete the square on the x's.

zepdrix (zepdrix):

Oops, a's instead of x's :)

zepdrix (zepdrix):

We currently have:\[\Large\rm 3\color{orangered}{a^2}+18\color{orangered}{a}+76=8\]Let's subtract 76 to the other side,\[\Large\rm 3\color{orangered}{a^2}+18\color{orangered}{a}=8-76\]\[\Large\rm 3\color{orangered}{a^2}+18\color{orangered}{a}=-68\]We need to get that leading coefficient off of the a^2 term. So we'll divide each side by 3,\[\Large\rm \color{orangered}{a^2}+\frac{18}{3}\color{orangered}{a}=-\frac{68}{3}\]

zepdrix (zepdrix):

What'dyou think Tiff? :d Following along?

zepdrix (zepdrix):

\[\Large\rm \color{orangered}{a^2+6a}=-\frac{68}{3}\]

OpenStudy (anonymous):

Yep!!

zepdrix (zepdrix):

So now we've got it in that nice form that we were looking for!\[\Large\rm a^2+ba=stuff\]To complete the square, we take half of our b coefficient and square it, \(\Large\rm \left(\dfrac{b}{2}\right)^2\) This will be the value that we'll add to each side. It will complete our square for us.

zepdrix (zepdrix):

So what is our b coefficient? What value are we going to end up adding to each side? What do you think tiff tiff? :O

OpenStudy (anonymous):

9!

zepdrix (zepdrix):

Ok great, our b coefficient is the 6, half of that is 3, squared is 9.\[\Large\rm \color{orangered}{a^2+6a}+9=-\frac{68}{3}+9\]

OpenStudy (anonymous):

would it be (a+3)^2=-41/3?

zepdrix (zepdrix):

Mmm yes very good! :) And depending on what you are trying to accomplish.. you would do one two things next. If you wanted to `solve for a`, you would take the square root of each side. Or more commonly, we're trying to often put something into `vertex form`, so we would want to add our 41/3 to the other side.

OpenStudy (anonymous):

i'm trying to find a

zepdrix (zepdrix):

Or were we just trying to complete the square on a? In that case I guess it wouldn't matter much where the 41/3 is, but it's nice to leave it with the other stuff,\[\Large\rm (a+3)^2+\frac{41}{3}=0\]

zepdrix (zepdrix):

If you're trying to `solve for a`, then we didn't want to go through the trouble of completing the square :) There's a much quicker way to solve for a. Unless that's the approach your teacher wanted you to take.

OpenStudy (anonymous):

can there be a square root on the denominator?

zepdrix (zepdrix):

\[\Large\rm \sqrt{(a+3)^2}=\sqrt{-\frac{41}{3}}=\frac{\sqrt{-41}}{\sqrt3}\]Yes.

OpenStudy (anonymous):

Wow! Okay, thank you sooooooo much!

zepdrix (zepdrix):

Think you can handle it from there? :) Remember how to deal with imaginary numbers and stuff? Bahh I gotta go get ready for work >.< No Problem \c:/ Good job.

OpenStudy (anonymous):

Yep, i remember! thank you again!